OpenStudy (anonymous):

Solve the system, using substitution. What is the n coordinate of the solution? 2m - 3n = 6 and m = 3 - 2n A. 2 B. 0 C. -4 D. 2 E. no solution

OpenStudy (p0sitr0n):

isolate either variable: take m since its already isolated. replace m into the first equation 2(3-2n)-3n=6 6-4n-3n=6 6-7n=6 -7n=6-6 n=0